OpenStudy (anonymous):

How do I find an equation of the line containing the given pair of points. (-7,-9) and (-5,-6)

OpenStudy (anonymous):

first slope: (-6+9)/(-5+7)=3/2 y=3/2(x) +b substitute one of the sets of points -6=-15/2+b b=1.5 y=3/2(x)+1.5